# Sample size and power calculation for propensity score analysis of observational studies

### v2

@@ 6 Real application: Right heart catheterization study
This section illustrates the proposed methods using simulations based on the right heart catheterization (RHC) observational study, a widely used benchmark dataset (Connors et al.,, 1996). Departing from the simulations in Section 5, here we do not impose any assumptions on the data generating model and thus provide a “fairer” assessment of the performance of our methods in real world settings. The RHC study was to evaluate the causal effect of the RHC procedure (binary treatment) among hospitalized critically ill patients. There are 5735 patients, each with 51 covariates (20 continuous, 25 binary, 6 categorical). The outcome is the binary survival status at 30 days after admission, with $Y=1$ meaning death. In total, 2184 ($38.1\%$) of the patients received RHC and the rest did not. The death rate is 38.0% and 30.6% in the treatment and control group, respectively. Using a logistic model with main effects of each covariate, we estimated the propensity scores and show their distributions by group in Appendix C.1 (Figure 6), suggesting a moderate degree of overlap. We numerically compute the true overlap coefficient defined in (10) to be 0.835. We also used the proposed normal approximation in Section 3.2, $W_{e}\sim\mathsf{N}(-0.701,2.189)$, and estimate the overlap coefficient to be 0.822, which is close to the true value.
The Hájek estimate of the ATE is 0.066. We use bootstrap to calculate the power of the two-sided hypothesis test (4) based on the Hájek estimator. Specifically, we generate $B=1000$ replicates, each with 5735 rows randomly drawn from the original data with replacement. We calculate the Hájek estimate of the ATE and the associated $95\%$ confidence interval in each replicate. Among the 1000 replicates, 983 intervals do not cover zero, giving an estimated power of 0.983 (95% CI: 0.975 to 0.991). Based on the RHC data, we compute the summary quantities $\theta=(r,\phi,E_{1},E_{0},S_{1},S_{0},R_{1},R_{0})$ as defined in Section 3 to be $(0.38,0.84,0.38,0.31,0.24,0.21,0.01,-0.02).$ The correlation $R_{z}=\mathrm{cor}[Y,W_{e}\mid Z=z]$ is close to zero, despite that the RHC study is not a randomized trial. This could be due to the fact that the outcome is binary and thus the correlation may be misleading.
Now consider the reversed setting of sample size calculation: we want to design an observational study to replicate the RHC study. We do not have the individual data but only some summary information, based on which we calculate the sample size required to achieve the power of 0.983 with the desired effect size being 0.066. Applying the proposed method, we find that the minimal sample size is 8349, which is 45.6% larger than the actual sample size 5735. For the power values at the two endpoints of the CI, the minimal sample sizes are 7706 and 9384, respectively. There are several potential reasons for the discrepancy: (i) we used the estimated instead of the true propensity score, (ii) we treated the binary outcome as normal, and (iii) the sample size for the ATE is sensitive to minor change in power which is so close to 1. In contrast, the sample size calculated by the two-sample $z$-test is 3426 (95% CI: 3163 to 3851), nearly half of the actual sample size, leading to a vastly underpowered study. Given that any real study would have data attrition due to missing data and extreme weights, a conservative (overpowered) estimate of the sample size is arguably more desirable than the opposite.
We repeat the process for the ATT and ATO. The Hájek estimate of the ATT is 0.070 with power 0.966 (95% CI: 0.955 to 0.977). To achieve the power of 0.966 (0.955 to 0.977) with an effect sample size 0.070, we calculate the minimal sample size to be 4538 (4233 to 4956), which is roughly 10% below the true sample size 5735. Unlike the ATE where we overestimate the sample size, our estimate for the ATT sample size is not guaranteed conservative. For the ATO, the CIs do not cover 0 in any of the replicates, indicating that the sample is overpowered for the ATO. Therefore, we subsample 3000 rows completely at random, with which the ATO is estimated to be 0.059. Among the 1000 replicates, 895 CIs do not cover 0, leading to an estimated power of 0.895 (95% CI: 0.876 to 0.914). Reversing the process, we calculate that the sample size to be 3401 (95% CI: 3197 to 3643), which is slightly above the true sample size $3000$.
@@
Figure 3 shows the power curves of the ATE given different values of the overlap coefficient $\phi$ and the corelation $R_{1}=R_{0}=R$ while fixing other summary inputs $(r,E_{1},E_{0},S_{1},S_{0})$ as those calculated from the RHC data. It shows that, to achieve the same power, the minimal sample size would increase as either the overlap decreases or the correlation $R$ increases. Both patterns are as expected from the theory. The SNRs for both treatment groups are around 0.29, indicating that the absolute correlation $|R|$ is bounded within $[0,0.46]$, which we set as the range in our power curves.
@@
Figure 3: Power and sample size curves of an emulated RHC study with an effect size $\tau=0.067$ under different values of the overlap coefficient $\phi$ (left) and the correlation $R=R_{1}=R_{0}$ (right).
@@ 7 Discussion
This paper formulates sample size and power calculations for the propensity score analysis of causal inference with observational data. We devise theoretically justified procedures and analytical formulas to identify the three key components for sample size calculations based on commonly available and interpretable summary statistics from previous studies. We elucidate the critical role of covariate overlap between treatment groups in determining the sample size and power. Our method is applicable to continuous and binary outcomes, and the general class of weighted average treatment effect estimands that includes ATE, ATT, ATO. We develop the associated R package PSpower.
An important extension is to time-to-event outcomes. In principle, the general structure of sample size calculations and in particular the identification of the propensity score distribution in Section 3 applies to all types of outcomes. However, censoring adds complexity to the identification of the potential outcome distributions. The hazard ratio is the dominant estimand for time-to-event outcomes. It is tied to the Cox proportional hazard outcome model and its causal interpretation stands only in certain settings (Hernán,, 2010). The existing literature for power calculation of multivariate Cox regression used the variance inflation approximation (Hsieh and Lavori,, 2000; Scosyrev and Glimm,, 2019), but it does not apply to causal analysis. A natural direction for sample size calculations of causal survival studies is to analyze the variance of the inverse probability weighted Cox model. Though closed-form solutions for the sample size calculation of the resulting marginal hazard ratio are unlikely to be available, easily computable approximations are expected to exist following the similar structure developed in this paper.
Appendix
Appendix A. Proof of Propositions and Theorems
A.1. Proof of Proposition 1
@@
$\Pr\left(\tau+\epsilon\sqrt{\frac{V_{0}}{N}}>z_{1-\alpha}\sqrt{\frac{V}{N}}\right)$ $=\Pr\left(\epsilon>z_{1-\alpha}\sqrt{\frac{V}{V_{0}}}-\tau\sqrt{\frac{N}{V_{0}}}\right)$ (26)
$=1-\Phi\left(\frac{z_{1-\alpha}\sqrt{V}-\tau\sqrt{N}}{\sqrt{V_{0}}}\right).$ (27)
@@
$N=\frac{\left(z_{1-\alpha}\sqrt{V}-z_{1-\beta}\sqrt{V_{0}}\right)^{2}}{\tau^{2}}.$ (28)
@@
$N\leq\frac{V}{\tau^{2}}\left(z_{1-\alpha}-z_{1-\beta}\right)^{2}=\frac{V}{\tau^{2}}\left(z_{1-\alpha}+z_{\beta}\right)^{2}.$ (29)
A.2. Special case of the Lyapunov Central Limit Theorem
@@
$\lim_{n\rightarrow\infty}\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left[|Z_{i}-\mu_{i}|^{2+\delta}\right]=0$ (30)
@@
$\frac{1}{s_{n}}\sum_{i=1}^{n}(Z_{i}-\mu_{i})\Rightarrow\mathsf{N}(0,1).$ (31)
@@ A.3. Proof of Proposition 2
$\mathbb{E}[Z]=\mathbb{E}[\mathbb{E}[Z\mid U]]=\mathbb{E}[U]=\frac{a}{a+b}.$ (32)
@@ A.4. Proof of Proposition 3
$h^{\prime}(k)=r\psi\left(kr+\frac{1}{2}\right)-r\psi(kr)+(1-r)\psi\left(k(1-r)+\frac{1}{2}\right)-(1-r)\psi(k(1-r))-\frac{1}{k}.$ (34)
@@
$\psi(x)-\psi(y)=\int_{0}^{1}\frac{t^{y-1}-t^{x-1}}{1-t}\,\mathrm{d}t,\quad\forall x,y>1,$ (35)
@@
$h^{\prime}(k)$ $=r\int_{0}^{1}\frac{t^{kr-1}\left(1-\sqrt{t}\right)}{1-t}\,\mathrm{d}t+(1-r)\int_{0}^{1}\frac{t^{k(1-r)-1}\left(1-\sqrt{t}\right)}{1-t}\,\mathrm{d}t-\frac{1}{k}$ (36)
$=\int_{0}^{1}\frac{rt^{kr-1}+(1-r)t^{k(1-r)-1}}{1+\sqrt{t}}\,\mathrm{d}t-\frac{1}{k}.$ (37)
@@
$h^{\prime}(k)$ $=\int_{0}^{1}\frac{2ru^{2kr}+2(1-r)u^{2k(1-r)}}{u(1+u)}\,\mathrm{d}u-\frac{1}{k}$ (38)
$=\frac{1}{k}\left\{\int_{0}^{1}\frac{\mathrm{d}\left(u^{2kr}+u^{2k(1-r)}\right)}{1+u}-1\right\}$ (39)
$=\frac{1}{k}\left\{\left.\frac{u^{2kr}+u^{2k(1-r)}}{1+u}\right|_{0}^{1}+\int_{0}^{1}\frac{u^{2kr}+u^{2k(1-r)}}{(1+u)^{2}}\,\mathrm{d}u-1\right\}$ (40)
$=\frac{1}{k}\int_{0}^{1}\frac{u^{2kr}+u^{2k(1-r)}}{(1+u)^{2}}\,\mathrm{d}u>0.$ (41)
A.5. Proof of Equation (14).
@@
$E_{z}$ $\equiv\mathbb{E}[Y\mid Z=z]=\mathbb{E}[a_{z}W_{e}+\epsilon_{z}\mid Z=z]=a_{z}\mathbb{E}[W_{e}\mid Z=z]+\mu_{z};$ (42)
$S_{z}$ $\equiv\mathbb{V}[Y\mid Z=z]=\mathbb{V}[a_{z}W_{e}+\epsilon_{z}\mid Z=z]=a_{z}^{2}\mathbb{V}[W_{e}\mid Z=z]+\sigma_{z}^{2};$ (43)
$R_{z}$ $\equiv\mathrm{cor}[Y,W_{e}\mid Z=z]=\mathrm{cor}[a_{z}W_{e}+\epsilon_{z},W_{e}\mid Z=z]=\frac{a_{z}\mathbb{V}[W_{e}\mid Z=z]}{\sqrt{S_{z}\mathbb{V}[W_{e}\mid Z=z]}}.$ (44)
By (44),
$a_{z}=\rho_{z}\sqrt{\frac{S_{z}}{\mathbb{V}[W_{e}\mid Z=z]}},$ (45)
which, plugged into (42) and (43), yields
$\mu_{z}=E_{z}-a_{z}\mathbb{E}[W_{e}\mid Z=z],\quad\sigma_{z}^{2}=(1-R_{z}^{2})S_{z}.$ (46)
A.6. Proof of the analytical form of variance $V$ in Theorem 1.
@@
$V=\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right],$ (47)
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=\mathbb{E}\left[\frac{a_{1}^{2}(W_{e}-\mu_{e})^{2}+\epsilon_{1}^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]=a_{1}^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (48)
@@
$\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (49)
$=\int\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (50)
$=\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u+\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (51)
$=1+\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (52)
@@
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (53)
$=\int u^{2}\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (54)
$=\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u+\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (55)
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (56)
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=a_{1}^{2}\sigma_{e}^{2}+\sigma_{1}^{2}+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (57)
@@
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=\mathbb{E}\left[\frac{a_{0}^{2}(W_{e}-\mu_{e})^{2}+\epsilon_{0}^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]=a_{0}^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{1}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (58)
@@
$\mathbb{E}\left[\frac{1}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (59)
$=\int\left(1+e^{u+\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (60)
$=\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}+\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (61)
$=1+\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (62)

### v3

@@ 4.2 Real data based simulations
This section illustrates the proposed methods using simulations based on a widely used benchmark data from the Right Heart Catheterization observational study (Connors et al.,, 1996), which was conducted to evaluate the causal effect of the diagnostic cardiological procedure of right heart catheterization (binary treatment) among hospitalized critically ill patients. The purpose of our simulation is not to re-design an existing study; instead, we want to assess the performance of our methods in a real world setting where we do not know the true data generating model, presenting a more realistic and challenging setting. The data contain 5735 patients, each with 51 covariates (20 continuous, 25 binary, 6 categorical). The outcome is the binary survival status at 30 days after admission, with $Y=1$ meaning death. In total, 2184 ($38.1\%$) of the patients received the treatment and the rest did not. The death rate is 38.0% and 30.6% in the treatment and control group, respectively.
The simulation proceeds as follows. We first calculate the summary parameters $(r,\phi,\rho^{2})$ and effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$ from the real data. We take these values as the only inputs of a hypothetical power analysis, without access to the whole data. We apply the proposed method with these inputs to compute the minimal sample size needed to achieve a certain level of power in detecting the pre-fixed effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$. We then reverse the process to estimate the power under a sample of that calculated size. The closer the estimated power is to the nominal power (e.g. 0.8), the more accurate the our method is. We did not compare with the method of Shook-Sa and Hudgens, (2022) because it requires the distribution of weights, which is assumed not available in this study.
Following the above outline, from the real data we estimate the propensity scores using a logistic model with the main effects of each covariate (the distributions by group are shown in Supplementary Material C.1 (Figure 6). Based on these estimated scores, we numerically compute the true overlap coefficient $\rho$ defined in (9) to be 0.835. We also used the proposed normal approximation in Section 3.2, $W_{e}\sim\mathsf{N}(-0.701,2.189)$, and estimate the overlap coefficient to be 0.822, which is close to the true value, showing the precision of our proposed Beta approximation. The Hájek estimate of the standardized effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$ is 0.14. The treatment proportion is $r=0.381$ and $\rho^{2}=0.000$. With these inputs alone, we apply formula (17) to calculate the minimal sample size to achieve 0.8 power to be 3625. We then reverse the process to calculate the actual power under a sample of 3625 units. Specifically, we generate $B=1000$ replicates, each with 3625 rows randomly drawn from the original data with replacement. We calculate the Hájek estimate of $\tau^{\mbox{\tiny{ATE}}}$ and the associated $95\%$ confidence interval in each replicate. Among the 1000 replicates, 924 intervals do not cover zero, giving an estimated power of 0.924 with 95% confidence interval (0.908 – 0.940), which is larger than the nominal power of 0.8. Potential reasons for the discrepancy include: (i) we used the estimated instead of the true propensity scores, and (ii) we treated the binary outcome as Gaussian. We also computed the sample size calculated by the two-sample $z$-test, which is 1567, leading to a vastly underpowered estimate with the power of only 0.55. Given that any real study would have data attrition due to missing data and extreme weights, a conservative (overpowered) estimate of the sample size is arguably more desirable than the opposite.
@@
Figure 3 shows the power curves of $\tau^{\mbox{\tiny{ATE}}}$ given different values of the overlap coefficient $\phi$ and the correlation $\rho_{1}^{2}=\rho_{0}^{2}=\rho^{2}$ while fixing other summary inputs as those calculated from the Right Heart Catheterization data. It shows that, to achieve the same power, the minimal sample size would increase as either the overlap decreases or the correlation $\rho$ increases. Both patterns are as expected from the theory. The $R^{2}$ value is 0.21, which serves as the upper bound of $\rho^{2}$.
We repeat the process for the average treatment effect for the treated and the overlap population, respectively. The results are relegated to Supplementary Material D.5.
@@
Figure 3: Power and sample size curves of an emulated Right Heart Catheterization study with an effect size $\widetilde{\tau}=0.14$ under different overlap coefficient $\phi$ (left) and the correlation $\rho=\rho_{1}=\rho_{0}$ (right).
@@ 5 Discussion
This paper investigates the theoretical foundation for power analysis in the propensity score analysis of causal inference with observational data. We show that, compared with the power calculation of randomized trials, users only need to specify two parameters for emulated trials, which quantifies the strength of the confounder-treatment and the confounder-outcome association, respectively. We propose interpretable and easily computable statistics to characterize these two parameters. Our method bypasses distributional assumptions on the multivariate covariates and generalizes the existing variance inflation factor method. We provide new theoretical insights on the sufficient summary information for power calculation in observational studies.
There are several important extensions of our method in observational studies, including repeated outcomes (Liu and Wu,, 2005), time-to-event outcomes (Hsieh and Lavori,, 2000), and multilevel data (Arpino and Mealli,, 2011). Propensity weighting estimators for causal effects are available in all these settings. Therefore, we can readily adapt the general structure of power calculation developed in this paper. For example, for studies with repeated outcomes, a common estimand is the time-averaged difference; we can analyze the variance of an inverse probability weighted marginal outcome model within-subject correlation between outcomes. For time-to-event outcomes, a common estimand is the marginal causal hazard ratio from the Cox proportional hazard model; we can analyze the variance of the inverse probability weighted Cox model.
Supplementary Materials
A Proof of Propositions and Theorems
A.1 Proof of (5)
@@
$\Pr\left(\tau+\epsilon\sqrt{\frac{V_{0}}{N}}>z_{1-\alpha}\sqrt{\frac{V}{N}}\right)$ $=\Pr\left(\epsilon>z_{1-\alpha}\sqrt{\frac{V}{V_{0}}}-\tau\sqrt{\frac{N}{V_{0}}}\right)$ (A.1)
$=1-\Phi\left(\frac{z_{1-\alpha}\sqrt{V}-\tau\sqrt{N}}{\sqrt{V_{0}}}\right).$ (A.2)
@@
$N=\frac{\left(z_{1-\alpha}\sqrt{V}-z_{1-\beta}\sqrt{V_{0}}\right)^{2}}{\tau^{2}}.$ (A.3)
@@
$N\leq\frac{V}{\tau^{2}}\left(z_{1-\alpha}-z_{1-\beta}\right)^{2}=\frac{V}{\tau^{2}}\left(z_{1-\alpha}+z_{\beta}\right)^{2}.$ (A.4)
A.2 Special case of the Lyapunov Central Limit Theorem
@@
$\lim_{n\rightarrow\infty}\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left[|Z_{i}-\mu_{i}|^{2+\delta}\right]=0$ (A.5)
@@
$\frac{1}{s_{n}}\sum_{i=1}^{n}(Z_{i}-\mu_{i})\Rightarrow\mathsf{N}(0,1).$ (A.6)
@@ A.3 Proof of Proposition 1
$\mathbb{E}[Z]=\mathbb{E}[\mathbb{E}[Z\mid U]]=\mathbb{E}[U]=\frac{a}{a+b}.$ (A.7)
@@ A.4 Proof of Proposition 2
$h^{\prime}(k)=r\psi\left(kr+\frac{1}{2}\right)-r\psi(kr)+(1-r)\psi\left(k(1-r)+\frac{1}{2}\right)-(1-r)\psi(k(1-r))-\frac{1}{k}.$ (A.9)
@@
$\psi(x)-\psi(y)=\int_{0}^{1}\frac{t^{y-1}-t^{x-1}}{1-t}\,\mathrm{d}t,\quad\forall x,y>1,$ (A.10)
@@
$h^{\prime}(k)$ $=r\int_{0}^{1}\frac{t^{kr-1}\left(1-\sqrt{t}\right)}{1-t}\,\mathrm{d}t+(1-r)\int_{0}^{1}\frac{t^{k(1-r)-1}\left(1-\sqrt{t}\right)}{1-t}\,\mathrm{d}t-\frac{1}{k}$ (A.11)
$=\int_{0}^{1}\frac{rt^{kr-1}+(1-r)t^{k(1-r)-1}}{1+\sqrt{t}}\,\mathrm{d}t-\frac{1}{k}.$ (A.12)
@@
$h^{\prime}(k)$ $=\int_{0}^{1}\frac{2ru^{2kr}+2(1-r)u^{2k(1-r)}}{u(1+u)}\,\mathrm{d}u-\frac{1}{k}$ (A.13)
$=\frac{1}{k}\left\{\int_{0}^{1}\frac{\mathrm{d}\left(u^{2kr}+u^{2k(1-r)}\right)}{1+u}-1\right\}$ (A.14)
$=\frac{1}{k}\left\{\left.\frac{u^{2kr}+u^{2k(1-r)}}{1+u}\right|_{0}^{1}+\int_{0}^{1}\frac{u^{2kr}+u^{2k(1-r)}}{(1+u)^{2}}\,\mathrm{d}u-1\right\}$ (A.15)
$=\frac{1}{k}\int_{0}^{1}\frac{u^{2kr}+u^{2k(1-r)}}{(1+u)^{2}}\,\mathrm{d}u>0.$ (A.16)
A.5 Proof of Equation (12).
@@
$E_{z}$ $\equiv\mathbb{E}[Y\mid Z=z]=\mathbb{E}[a_{z}W_{e}+\epsilon_{z}\mid Z=z]=a_{z}\mathbb{E}[W_{e}\mid Z=z]+\mu_{z};$ (A.17)
$S_{z}^{2}$ $\equiv\mathbb{V}[Y\mid Z=z]=\mathbb{V}[a_{z}W_{e}+\epsilon_{z}\mid Z=z]=a_{z}^{2}\mathbb{V}[W_{e}\mid Z=z]+\sigma_{z}^{2};$ (A.18)
$\rho_{z}$ $\equiv\mathrm{cor}[Y,W_{e}\mid Z=z]=\mathrm{cor}[a_{z}W_{e}+\epsilon_{z},W_{e}\mid Z=z]=\frac{a_{z}\mathbb{V}[W_{e}\mid Z=z]}{\sqrt{S_{z}\mathbb{V}[W_{e}\mid Z=z]}}.$ (A.19)
By (A.19),
$a_{z}=\rho_{z}\sqrt{\frac{S_{z}^{2}}{\mathbb{V}[W_{e}\mid Z=z]}},$ (A.20)
which, plugged into (A.17) and (A.18), yields
$\mu_{z}=E_{z}-a_{z}\mathbb{E}[W_{e}\mid Z=z],\quad\sigma_{z}^{2}=(1-\rho_{z}^{2})S_{z}^{2}.$ (A.21)
A.6 Proof of the analytical form of variance $V$ in Theorem 2.
@@
$V=\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right],$ (A.22)
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=\mathbb{E}\left[\frac{a_{1}^{2}(W_{e}-\mu_{e})^{2}+\epsilon_{1}^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]=a_{1}^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (A.23)
@@
$\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.24)
$=\int\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.25)
$=\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u+\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.26)
$=1+\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (A.27)
@@
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.28)
$=\int u^{2}\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.29)
$=\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u+\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.30)
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.31)
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=a_{1}^{2}\sigma_{e}^{2}+\sigma_{1}^{2}+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.32)
@@
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=\mathbb{E}\left[\frac{a_{0}^{2}(W_{e}-\mu_{e})^{2}+\epsilon_{0}^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]=a_{0}^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{1}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (A.33)
@@
$\mathbb{E}\left[\frac{1}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.34)
$=\int\left(1+e^{u+\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.35)
$=\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}+\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.36)
$=1+\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (A.37)